P(z>0.23 or z<-0.23) = P(-3<z<O) = P(-0.23<z<2.35) = P(z>1.47 or z<-1.47) = P(z<-2 or z>-1) = P(-3<z<3) -
Answer #1

Ans. We calculate all probabilities using R:-

1. P(z>0.23 or z<-0.23) =P(z>0.23 ) + P( z<-0.23)

> pnorm(.23,0,1,lower.tail=F)+pnorm(-0.23,0,1)
[1] 0.8180918

P(z>0.23 or z<-0.23) = 0.8181

2. P(-3<z<0) = P(z<0) - P(z<-3)

> pnorm(0,0,1)-pnorm(-3,0,1)
[1] 0.4986501

P(-3<z<0) = 0.4987

3. P(-0.23<z<2.35) = P(z<2.35) - P(z<-0.23)

> pnorm(2.35,0,1)-pnorm(-0.23,0,1)
[1] 0.5815674

P(-0.23<z<2.35) = 0.5816

4. P(z>1.47 or z<-1.47) =P(z>1.47 ) + P( z<-1.47)

> pnorm(1.47,0,1,lower.tail=F)+pnorm(-1.47,0,1)
[1] 0.1415618

P(z>1.47 or z<-1.47) = 0.1416

5. P(z<-2 or z>-1) = P(z<-2) + P(z>-1)

> pnorm(-2,0,1)+pnorm(-1,0,1,lower.tail=F)
[1] 0.8640949

P(z<-2 or z>-1) = 0.8641

6. P(-3<z<3) = P(z<3) - P(z<-3)

> pnorm(3,0,1)-pnorm(-3,0,1)
[1] 0.9973002

P(-3<z<3) = 0.9973